Cot 7/3 pi, where 7/3 pi is radina measure, has the value of
Answer by stanbon(75887)

Cot[(7/3)pi], where 7/3 pi is radina measure, has the value of) ?
(7/3)pi = (2pi + (1/3)pi)
So, Cot[(7/3)pi] = Cot[(1/3)pi] = sqrt(3)
=============================================
Cheers,
Stan H.